# MIRROR: Multi-Modal Pathological Self-Supervised Representation Learning via Modality Alignment and Retention

### v1

@@ V Conclusion
In this paper, we introduced MIRROR, a novel multi-modal pathological SSL framework designed for joint pre-training of histopathology and transcriptomics data. Our approach is tailored to align modality-shared information while retaining modality-specific unique features. Additionally, it employs a style clustering module to reduce redundancy and preserve disease-relevant representations. Furthermore, we proposed a novel transcriptomics data preprocessing pipeline to efficiently identify disease-related genes, resulting in refined, disease-focused transcriptomics datasets. The proposed method was trained and evaluated on two distinct cohorts from the TCGA dataset, demonstrating superior performance and underscoring the efficacy of our design. Future work includes expanding evaluations to additional tasks and cohorts, as well as developing more advanced approaches for extracting information from raw transcriptomics data.

### v2

@@ V Conclusion
In this paper, we introduced MIRROR, a novel multi-modal pathological SSL framework designed for joint pre-training of histopathology and transcriptomics data. Our approach is tailored to align modality-shared information while retaining modality-specific unique features. Additionally, it employs a style clustering module to reduce redundancy and preserve disease-relevant representations. Furthermore, we proposed a novel transcriptomics data preprocessing pipeline to efficiently identify disease-related genes, resulting in refined, disease-focused transcriptomics datasets. The proposed method was trained and evaluated on four distinct cohorts from the TCGA dataset, demonstrating superior performance and underscoring the efficacy of our design. Future work includes expanding evaluations to additional tasks and cohorts, as well as developing more advanced approaches for extracting information from raw transcriptomics data.